OpenStudy (anonymous):

HELP Satellite. Cubed root of 512x^2y^4/8x^5y. Simplify the expression and express the answer using rational exponents. Assume that x and y denote positive numbers.

OpenStudy (anonymous):

\[\frac{\left(512x^2y^2\right)}{8x^{5y}}\] is that it?

OpenStudy (anonymous):

or \[\frac{\left(512x^2y^2\right)}{8x^5y}\]

OpenStudy (anonymous):

divide the numbers and subtract the exponents to get \[64x^{2-5}y^{2-1}=64x^{-3}y\]

OpenStudy (anonymous):

if you are told to use "positive exponents only" then rewrite as \[\frac{64y}{x^3}\] because \[x^{-3}=\frac{1}{x^3}\]